OpenStudy (anonymous):

A merchant has 5 pounds of mixed nuts that cost $30. He wants to add peanuts that cost $1.50 per pound and cashews that cost $4.50 per pound to obtain 50 pounds of a mixture that costs $2.90 per pound. How many pounds of peanuts are needed?

OpenStudy (amistre64):

a + b = t ax + by = tz -ax -bx = -tx ax +by = tz ------------- b(y-x) = t(z-x) b@y = t(z-x)/(y-x)

OpenStudy (amistre64):

5 pounds of mixed nuts $30. ; what is this for? (b)peanuts (y)$1.50 per pound cashews that cost (x)$4.50 per pound to obtain (t)50 pounds (z)$2.90 per pound. How many pounds of (b)peanuts are needed? b@y = 50(2.90-4.50)/(1.50-4.50)

OpenStudy (amistre64):

looks to be about 26 and 2/3

OpenStudy (anonymous):

ugh see im so confused how you got the equations

OpenStudy (amistre64):

80/3 (1.5) + 70/3(4.5) = 50(2.9) 145 = 145

OpenStudy (amistre64):

26 and 2/3 lbs of peanuts

OpenStudy (amistre64):

the equations are standard for these types of problems; knowing what to fill in to the spots is the challange

OpenStudy (anonymous):

see i dont understand :( my teacher didnt explain it well and our whole class is lost

OpenStudy (amistre64):

you know you need 2 different amounts; and they have to add up to a (t)otal of 50 lbs: a + b = t ; you have prices to account for; a costs x much; b costs y much and when it is all said and done the (t)otal needs to cost z much

OpenStudy (amistre64):

a + b = t ax + by = tz this is a system of eqatuons where a and b have to be the same amount and we can find them quite readily

OpenStudy (whpalmer4):

End result is 50 lbs of nuts at $2.90/lb = $145 total cost. We've already got 5 lbs of nuts that cost $30, so the remaining 45 lbs must cost $145-$30 = $115. That works out to an average cost of $115/45 per pound. $1.50 * p + $4.50 * (1-p) = $115/45 (p being the fraction of a pound of the added nuts mixture that is peanuts, and 1-p being the fraction of a pound that is cashews). If we solve that equation for p, we get p = 0.648. Multiply that fraction by 45 pounds, and we see that we need 29.17 lbs of peanuts. Checking our work, 29.17*1.5 + (45-29.17)*4.5 = $115.

OpenStudy (amistre64):

we can either graph the lines the equation will produce; or i just eliminate a variable and solve for the remainingone

OpenStudy (amistre64):

is that what that first part was saying? hunh ,,,,,,

OpenStudy (amistre64):

i couldnt figure out what the 5lbs at 30 had to do with anything :)

OpenStudy (amistre64):

ida feed it to the hogs and made a new batch

OpenStudy (anonymous):

ugh i just give up none of this is making sense

OpenStudy (anonymous):

thanks it kind of makes sense now